Cable TV choices: A telecommunications company provided its cable TV subscribers with free access to a new sports channel for a period of one month. It then chose a sample of 412 televisionviewers and asked them whether they would be willing to pay. The marketing director of the company claims that the percentage of all of its subscribers who would pay for the channel differs from 6%. Can you conclude that the director's claim is true? Use alpha=0.01 level of significance and the critical value method.

Part B question Find the critical values. Round your answer to at least three decimal places. Critical values are ________ and _________.
Answer #1

Since n * p * (1 - p) = 412 * 0.06 * 0.94 = 23.23, which is > 10, we use the z distribution.

The 2 tailed p values for = 0.01 are -2.576 and +2.576

_____________________________________

The rejection rule using these critical values is Reject H0, if Z test is > 2.576 or if Z test is < -2.576

___________________________________

The Test Statistic: = 33 / 412 = 0.08

The Decision: Since 1.72 lies in between -2.576 and 2.576, We Fail to Reject H0

The Conclusion: There isn't sufficient evidence at the 99% significance level to conclude that the percentage of all subscribers who would pay for the channel differs from 6%